# Dark Photon Oscillations in Waveguide

### v1

@@ III Expected Sensitivity
The illustrative square pulse ( shown in panel c)) is a carrier at $\omega_{0}$ modulated by a slow envelope $\Omega\ll\omega_{0}$ and decomposes into the dominant tone $\omega_{0}$, lower harmonics $\{\omega_{0}-n\Omega\}$, and higher harmonics $\{\omega_{0}-n\Omega\}$. This decomposition motivates the “D+H” strategy: when $\omega_{0}$ is tuned near the resonant mode the lower harmonics become evanescent and are filtered out by the guide, leaving the dominant tone plus higher harmonics (shown in panel d)) as the effective launched field. Note that there are some distortions of the waveform after a long propagation in the fiber, which will be shown in the supplemental material. Each spectral component propagates with phase velocity $v_{p,B_{i}}^{n}=\left(\omega_{0}+n\Omega\right)/k_{z,B_{i}}$. So the two eigenstates accumulate a component–dependent phase difference
@@
$\overline{\Delta\varphi_{0}}=\frac{4\varepsilon L}{\lambda}\int_{0}^{\infty}\frac{1}{\sqrt{x\zeta+2\alpha}+\sqrt{x\zeta}}\frac{1}{1+x^{2}}dx,$ (5)
@@
In all, the sensitivity in the OTDR is set by the fit uncertainty per measurement point, the coherence (linewidth) of the laser, and the effective baseline $L_{\rm eff}$: long, coherent fibers and narrow laser linewidths increase accumulated phase difference and hence reach, whereas finite coherence and practical noise set the achievable limit on the kinetic-mixing parameter. This approach therefore converts distributed, laboratory-scale optical sensing into a quantitative bound on dark-photon mixing without accessing the fiber interior or using auxiliary fields, while requiring only standard fiber-optic instrumentation and mode control. Anyway, we can measure backscatter-derived power $\hat{P}(z_{i})$ at a set of lengths ${\cal M}={\hat{P}(z_{1}),\ldots,\hat{P}(z_{n})}$. Then we can compute normalized values $P(z_{i})=\hat{P}(z_{i})e^{2\beta_{\xi}z_{i}}$ where $\beta_{\xi}$ is the measured attenuation coefficient, which removes monotonic loss so that residual spatial structure reveals oscillation. Fit $P(z)$ to a cosine model (or sinusoidal envelope appropriate to the two-eigenstate interference). A significant cosine component with the expected spatial period is evidence for oscillation. When $\lambda_{\rm osc}\gg L$ expand the cosine and use three-point measurements $P(0),P(L/2),P(L)$ to estimate $\Delta k_{z}L$ via newly defined quantity ${\cal E}(L)$
@@ V Acknowledgement
I Maxwell’s Equations and Mode Quantization in a Waveguide
@@ II The Propagation and Oscillation of Dark Photons
At exact resonance ($m_{\xi}=m_{A^{\prime}}$), $\theta\approx\pi/4$, so $\eta_{1}=\eta_{2}=1/\sqrt{2}$ and
@@
III Propagation of Pulsed Signals: Fourier Analysis
@@
$=$ $\frac{4\varepsilon L}{\lambda}\int_{0}^{\infty}\frac{1}{\sqrt{x\zeta+2\alpha}+\sqrt{x\zeta}}\frac{1}{1+x^{2}}dx\,.$
@@
Calculation of the Generally speaking, the experimental measurements of the oscillation is simple. Both power flow which is measured by a power meter, and the amplitude which is shown by the oscilloscope, can be chosen for the display in the panel a) of Fig. 2. Suppose the power flow are measured at a series of lengths for the prepared fiber with a radius $r_{i}$
@@
Then one may worry about that the mixing parameters $\eta_{1},\penalty 10000\ \eta_{2}$ are unknown, This can be easily steered clear of by analyzing the power flow at two different points, say $z=L/2,L$. The phase difference can be easily derived form the by the expansion Eq. (S.43)
$\frac{P(L/2;\penalty 10000\ r_{r})-P(0;\penalty 10000\ r_{r})}{P(L;\penalty 10000\ r_{r})-P(0;\penalty 10000\ r_{r})}\simeq\frac{1}{4}\left[1+\frac{1}{16}(\Delta k_{z}L)^{2}\right]\,.$ (S.49)
A much more precise $\Delta k_{z}L$ can be extracted by this relation if the oscillation wavelength $\lambda_{\text{osc}}$ is too much long. Three measured samples are enough for the fitting for the oscillation.

### v2

@@ III Expected Sensitivity
The illustrative square pulse ( shown in panel c)) is a carrier at $\omega_{0}$ modulated by a slow envelope $\Omega\ll\omega_{0}$ and decomposes into the dominant tone $\omega_{0}$, lower harmonics $\{\omega_{0}-n\Omega\}$, and higher harmonics $\{\omega_{0}+n\Omega\}$. This decomposition motivates the “D+H” strategy: when $\omega_{0}$ is tuned near the resonant mode the lower harmonics become evanescent and are filtered out by the guide, leaving the dominant tone plus higher harmonics (shown in panel d)) as the effective launched field. Note that there are some distortions of the waveform after a long propagation in the fiber, which will be shown in the supplemental material. Each spectral component propagates with phase velocity $v_{p,B_{i}}^{n}=\left(\omega_{0}+n\Omega\right)/k_{z,B_{i}}$. So the two eigenstates accumulate a component–dependent phase difference
@@
$\overline{\Delta\varphi_{0}}=\frac{4\varepsilon L}{\lambda}\int_{0}^{\infty}\frac{1}{\sqrt{x\zeta+2\varepsilon}+\sqrt{x\zeta}}\frac{1}{1+x^{2}}dx,$ (5)
@@
In all, the sensitivity in the OTDR is set by the fit uncertainty per measurement point, the coherence (linewidth) of the laser, and the effective baseline $L_{\rm eff}$: long, coherent fibers and narrow laser linewidths increase accumulated phase difference and hence reach, whereas finite coherence and practical noise set the achievable limit on the kinetic-mixing parameter. This approach therefore converts distributed, laboratory-scale optical sensing into a quantitative bound on dark-photon mixing without accessing the fiber interior or using auxiliary fields, while requiring only standard fiber-optic instrumentation and mode control. In the experiment, we can measure backscatter-derived power $\hat{P}(z_{i})$ at a set of lengths ${\cal M}={\hat{P}(z_{1}),\ldots,\hat{P}(z_{n})}$. Then we can compute normalized values $P(z_{i})=\hat{P}(z_{i})e^{2\beta_{\xi}z_{i}}$ where $\beta_{\xi}$ is the measured attenuation coefficient, which removes monotonic loss so that residual spatial structure reveals oscillation. Fit $P(z)$ to a cosine model (or sinusoidal envelope appropriate to the two-eigenstate interference). A significant cosine component with the expected spatial period is evidence for oscillation. When $\lambda_{\rm osc}\gg L$ expand the cosine and use three-point measurements $P(0),P(L/2),P(L)$ to estimate $\Delta k_{z}L$ via newly defined quantity ${\cal E}(L)$
@@ V Acknowledgement
I Maxwell’s Equations and Propagation Modes in a Waveguide
@@ II The Propagation and Oscillation of Dark Photons
At exact resonance ($m_{\xi}=m_{A^{\prime}}$), $\theta=\pi/4$, so $\eta_{1}=\eta_{2}=1/\sqrt{2}$ and
@@
III The Fourier analysis of Pulsed Signals and data processing for measurement
@@
$=$ $\frac{4\varepsilon L}{\lambda}\int_{0}^{\infty}\frac{1}{\sqrt{x\zeta+2\varepsilon}+\sqrt{x\zeta}}\frac{1}{1+x^{2}}dx\,.$
@@
Generally speaking, the experimental measurements of the oscillation is simple. Both power flow which is measured by a power meter, and the amplitude which is shown by the oscilloscope, can be chosen for the display in the panel a) of Fig. 2. Suppose the power flow are measured at a series of lengths for the prepared fiber with a radius $r_{i}$
@@
Then one may worry about that the mixing parameters $\eta_{1},\penalty 10000\ \eta_{2}$ are unknown, This can be easily steered clear of by analyzing the power flow at two different points, say $z=L/2,\penalty 10000\ L$. The phase difference can be easily derived form the expansion Eq. (S.43), and we can define a measurable quantity ${\cal E}(L;\penalty 10000\ r_{r})$
${\cal E}(L;\penalty 10000\ r_{r})=\frac{P(L/2;\penalty 10000\ r_{r})-P(0;\penalty 10000\ r_{r})}{P(L;\penalty 10000\ r_{r})-P(0;\penalty 10000\ r_{r})}\simeq\frac{1}{4}\left[1+\frac{1}{16}(\Delta k_{z}L)^{2}\right]\,.$ (S.49)
Take $\Delta k_{z}L=0.01\times 2\pi$ for the demonstration, then the measured $({\cal E}-0.25)\times 100\simeq 0.006$. A much more precise $\Delta k_{z}L$ can be extracted by this relation if the oscillation wavelength $\lambda_{\text{osc}}$ is too much long. Three measured samples are enough for the fitting for the oscillation.